# GT-SVQ: A Linear-Time Graph Transformer for Node Classification Using Spiking Vector Quantization

### v1

@@ 3.3 Overall Framework
where $L$ is the number of layers. We choose a simple fully connected layer as the classification head. It has been observed that projection blocks consisting of Multilayer Perceptrons (MLPs) and normalization layers exacerbate the overfitting problem on large-scale graphs for vanilla Transformers. Therefore, we discard redundant projection layers and retain only the self-attention module and the skip-connection structure He et al. (2016).
4 Experiments
4.1 Comparison with Existing Models
In this section, we conduct the experimental evaluation to show the effectiveness of GT-SVQ on 9 node classification datasets. A head-to-head comparison is conducted with state-of-the-art GNNs and GTs, based on their architectures. As shown in Table 1, these baselines fall into three categories: spike-based methods, Graph Transformer framework and vector quantization-based methods. The hyperparameters search strategy is deployed on both GT-SVQ and other baselines to get the optimal combinations of parameters. We perform all models on each dataset 5 times with different random seeds to report the mean and standard deviation. All experiments are conducted on a single NVIDIA RTX 4090 GPU. More experimental settings are presented in Appendix E.
@@ 4.1 Comparison with Existing Models
Model Components
SP GT VQ
SpikingGCNZhu et al. (2022) - -
SpikeNetLi et al. (2023a) - -
SpikeGCLLi et al. (2023b) - -
SpikeGTSun et al. (2024) -
NAGphormerChen et al. (2022) - -
NodeFormerWu et al. (2022) - -
SGFormerWu et al. (2024) - -
GOATKong et al. (2023) -
VQGraphYang et al. (2024) - -
GT-SVQ
Overall performance.
The experimental results are demonstrated in Table 2. As shown in the table, our method achieves competitive performance on all datasets, which is a significant advancement considering the information loss caused by low-precision spike count embeddings. Furthermore, GT-SVQ achieves predictive performances on par or even better than high-precision GT baselines. Compared to advanced linear-time GTs, our method shows slight improvements on 7 out of 9 datasets. The results indicate that spiking vector quantization dynamically generates the codebook via the fusion of SNNs and message propagation, and it effectively injects the graph inductive bias into the self-attention block to capture long-range node information. GT-SVQ outperforms other spike-based baselines on all datasets, which achieves an average improvement of 3.9%. Constructing the input sequence from propagation steps rather than repeating graphs multiple times endows GT-SVQ with desirable scalability. We also observe that GT-SVQ has comparable performances on two heterophily datasets (particularly Actor), and more discussions about heterophily are provided in Appendix D.
@@
#$\#nodes$ 2,708 3,327 19,717 18,333 34,493 7,600 28,281 169,343 2,449,029
#$\#edges$ 10,556 9,104 88,648 163,788 495,924 30,019 185,504 1,166,243 61,859,140
@@
GT-SVQ 84.7±0.8 74.0±0.5 80.6±0.4 93.7±0.4 96.2±0.0 39.1±0.2 65.7±0.1 72.4±0.3 74.8±0.4
4.2 Characteristics of Spiking Vector Quantization
To elaborately analyze the spiking vector quantization, we conduct a series of experiments on the GT-SVQ. We track the metrics including Codebook Usage and Accuracy to explore the following questions: (i) How does the size of latent embedding space affect our model? (ii) Is the spiking vector quantization a more efficient VQ alternative?
The influence of $|\tilde{\mathcal{C}}|$.
As aforementioned above, the number of propagation steps $T$ and the random feature dimension $D$ determine the size of latent space. In Figure 3, we evaluate the performances of GT-SVQ on Cora and CS datasets under different combinations between $D$ and $T$. The visualization results show that an excessively small latent space size will severely impair the performances of GT-SVQ. It generally improves accuracy as the latent space size increases. Essentially, $|\tilde{\mathcal{C}}|$ determines the upper bound of information the reconstructed codebook can store. When $|\tilde{\mathcal{C}}|$ is much smaller than the number of nodes, indiscriminately mapping a large number of nodes to the same codeword will bring significant information loss. Besides, we observe the diminishing gains from latent space size scaling, especially on small-scale graphs. We believe that node embeddings merely are transformed into low-precision counterparts rather than being compressed appropriately when feature dimensionality $D$ is too large. The quantization error introduced by spiking neurons will cause performance degradation. The propagation step $T$ correlates with the complexity of spiking patterns. As depicted in Figure 3, it suggests that $T>6$ may lead to subpar performance.
@@ 4.2 Characteristics of Spiking Vector Quantization
For exploring the codebook collapse problem in VQ-based graph models, we track the fraction of codewords that are used at least once after the training phase. The codebook usages between GT-SVQ and the other VQ-based graph methods are shown in Figure 4. More statistical results are attached in Appendix B. In previous works, the entire discrete latent space is represented as a pre-defined codebook. We select four combinations of $D$ and $T$ (($T=3/D=4$, $T=3/D=5$, $T=3/D=6$, $T=3/D=7$)) to match the latent space size ($2^{8}$, $2^{10}$, $2^{12}$, $2^{14}$). The results demonstrate that those methods using pre-defined codebooks suffer from the serious issue of codebook collapse. As the pre-defined codebook size increases, the codebook usage decreases. For GOAT, the average codebook usages are 10.6% and 8.7% on Cora and CS datasets. The codebook usages in VQGraph are slightly higher, which achieve 16.9% and 29.0%. When the latent space size exceeds $2^{10}$, the usages of both methods drop below 50%. It reveals that these VQ-based graph learning methods struggle with efficiently using large pre-defined codebooks. SVQ provides a generative solution, which dynamically selects the subset from a discrete spike count embedding space to reconstruct the codebook. It brings 100% codebook usage. Although the number of used codewords in GT-SVQ will be slightly larger than that of vanilla VQ counterparts in some cases, it is still significantly smaller than the size of pre-defined codebooks. In a nutshell, SVQ efficiently achieves better codebook utilization by implementing vector quantization from the perspective of SNNs.
Figure 3: Effects of random feature dimensionality $D$ and propagation step $T$ on Cora and CS.
Figure 4: Codebook usage under four different latent space sizes on Cora and CS. The solid-colored bar represents the number of used codewords, while the hatched bar denotes the size of the pre-defined codebook.
Datasets Metrics NAGphormer GOAT NodeFormer SGFormer SpikeGT Avg GT-SVQ
@@
4.3 Energy Efficiency Analysis
To verify the efficiency of GT-SVQ, we conduct energy efficiency analysis on CS, Physics, ogbn-arxiv and ogbn-products datasets based on the following metrics: the maximum memory usage, inference latency and theoretical energy consumption. The results are shown in Table 3. Besides, we also provide the theoretical energy consumption in the formula and other experimental details in Appendix C.
@@ 4.3 Energy Efficiency Analysis
The results show that GT-SVQ achieves the lowest inference latency across all datasets. Compared to other GTs, GT-SVQ with better performance achieves up to 130x lower inference latency. Most VQ-based methods tend to retrieve and replace node representations with the closest codewords during the inference phase. In SVQ, trained spiking neurons directly convert input features into codewords, which reduces the computational cost brought by retrieval. Integrating the reconstructed codebook in SVQ with linear-time attention modules brings a significant improvement in inference speed. For large-scale graphs with high feature dimensionality, repeating the graph to generate sequential inputs leads to high energy consumption. The results show that the previous spike-based GT consists of spike-driven Transformers and full-precision MPNNs, SpikeGT is hard to manifest its superiority in energy consumption on CS and Physics. Through generating low-dimensional random features and collecting sequential embeddings from propagation steps, GT-SVQ outperforms another spike-based GT across all datasets with acceptable theoretical energy consumption.
4.4 Ablation Study
@@
In this section, we conduct ablation studies to explore the influences of different components on predictive performances. To this end, we implement two classic linear-time attention modules (Performer Choromanski et al. (2020) and Linformer Wang et al. (2020)), two common spiking neurons (IF and LIF), two normalization algorithms (LayerNorm Ba et al. (2016) and STFNorm Xu et al. (2021)) and the vanilla VQ module to construct 7 GT-SVQ variants.
@@
The experimental results are demonstrated in Table 4. Although incorporating extra positional encodings enables Performer and Linformer to handle graph prediction tasks, they struggle to achieve good predictive performance on large-scale graphs like ogbn-arxiv. In GT-SVQ, the CGSA actively introduces the global structure information during attention score calculation. It suggests that developing graph structure-aware Transformers is a promising direction for scaling GTs on large-scale graphs. The choice of spiking neurons also affects the predictive performances of GT-SVQ. PLIF models with learnable membrane time constants and synaptic weights achieve slightly better performances in most cases. These neurons effectively improve the flexibility of SVQ. In addition, the well-designed normalization algorithm for spiking neurons, STFNorm outperforms the LayerNorm algorithm across all datasets. For spiking graph neural networks, the distribution of spiking node representations and corresponding normalization algorithms lack further exploration. We leave the designs of specific spiking neurons and normalization layers on the graph data for future work.
Table 4: Ablation studies on Pubmed, CS, Physics and ogbn-arxiv datasets. And $+x$ means replacing the original component in GT-SVQ with $x$.
@@
GT-SVQ 80.6±0.5 93.7±0.4 96.2±0.0 72.4±0.3
5 Conclusion
@@
In this study, we propose GT-SVQ, a linear-time Graph Transformer via spiking vector quantization. We find that message propagation patterns of different nodes can be encoded into the same codewords. Inspired by the observation, GT-SVQ not only considers spiking neurons as the low-power units, but also incorporates SNNs as learnable quantizers into Graph Transformers. It enables GT-SVQ to achieve faster inference speed and better predictive performance. Spike vector quantization which dynamically creates the spike count-based codebook provides a spiking perspective to address issues in current VQ-based graph learning methods. We believe that our work holds great promise from a neuroscientific perspective, and we hope it will inspire further research into more efficient Graph Transformers.
Appendix A Visualization Results of SVQ
To better demonstrate our observations, we perform SVQ and a simple message propagation model on the KarateClub dataset. Specifically, for the full-precision message propagation model, the intermediate embeddings across multi-step message propagation are fed into a mean aggregator to generate the node embeddings. The initial node features are sampled from a uniform distribution over the interval $(0,1)$, and the number of propagation steps is set to 4. For SVQ, the propagation steps are set to 10 and 4 to exhibit the impact of different spiking neuron settings on spike count embeddings. In the preliminary experiments, we observe that as the number of propagation steps increases, the neighborhood message embeddings of nodes in the same class become increasingly similar. It implies that we can generate the same representation for different nodes by capturing the dynamics in the propagation process. Benefiting from the powerful coding mechanism of SNNs for sequential data, we fed the above intermediate embeddings into spiking neurons to generate node representations based on spike counts. different nodes are represented by the same codeword, which means high-precision node embeddings can be encoded into a finite set of codewords from narrower and discrete spike count embedding space.
Furthermore, we perform the SVQ and the full-precision message propagation model on Cora, Citeseer and Pubmed datasets. In the implementation, the random features will serve as the initial membrane potential of spiking neurons. SVQ updates node representation by alternating propagation and normalization operations. The adjacency matrix with self-loops and $l_{2}$ normalization are selected as the propagation operator and normalization function, respectively. Visualization results are shown in Figure 5. The high-precision node representations (the leftmost plot in each line) can be projected into the finite set of low-precision codewords (the two rightmost plots in each line). Considering iteratively propagated messages as input currents of spiking neurons will generate expressive low-precision node vectors. Additionally, the results reveal the precision of codewords is governed by spiking neurons with different configurations. Higher threshold potentials always correspond to lower fire rates or spike counts, which indirectly drives SVQ to generate node representation with lower precision.
(a) The spike count visualization on Cora.
(b) The spike count visualization on Citeseer.
(c) The spike count visualization on Pubmed.
Figure 5: The visualization results of node representations which are decoded in the form of the spike count. The feature dimension is set to 16 and nodes are sorted by their categories. Brighter spots denote higher spike counts.
Appendix B Codebook Usage
In Table 5, we show the codebook usages and predictive performances of GT-SVQ and VQ-based graph learning baselines on Cora and CS. For VQGraph and GOAT, as the latent space size increases, the number of used codewords continues growing but the codebook usage decreases. GOAT is more sensitive to the latent space size. When the space size is excessively small, GOAT struggles with quantizing the learned node embedding vectors. For the latent space of size $2^{8}$, GOAT exhibits obvious performance gaps compared with the other two methods on both datasets. In GT-SVQ, the size of the reconstructed codebook $B$ is equal to the number of used codewords. We observe that enlarging the latent space by choosing a larger feature dimensionality $D$ will significantly increase the reconstructed codebook size. Therefore, the feature dimensionality should be decided according to the scale of graph data. As mentioned in Section 3.2, the complexity of CGSA is $O(NBd_{v})$. In practice, $B$ may be smaller than the dimensionality of intermediate embeddings. This contributes to the superior inference speed and computational efficiency of CGSA compared to some existing linear-time attention modules.
Table 5: Codebook analysis on Cora and CS datasets. For each, we compare GT-SVQ with the other two VQ-based graph models, VQGraph and GOAT. Two metrics are tracked, codebook usage (Usage) and accuracy (ACC).
Models Latent space size Cora CS
[…]
VQGraph $2^{8}$ 112/256   (43.8) 81.8±1.2 168/256   (65.6) 91.9±0.1
$2^{10}$ 172/1024   (16.8) 81.0±1.0 338/1024   (33.0) 92.0±0.5
$2^{12}$ 206/4096   (5.0) 81.4±1.1 542/4096   (13.2) 91.8±0.4
$2^{14}$ 317/16384   (1.9) 81.8±1.9 697/16384   (4.3) 91.7±0.4
GOAT $2^{8}$ 76/256   (29.7) 68.4±0.6 49/256   (19.1) 56.1±3.2
$2^{10}$ 99/1024   (9.7) 68.3±0.9 100/1024   (9.8) 92.2±0.3
$2^{12}$ 100/4096   (2.4) 72.2±0.4 154/4096   (3.8) 92.4±1.0
$2^{14}$ 116/16384   (0.7) 78.6±0.5 333/16384   (2.0) 92.0±0.8
GT-SVQ T=3,D=4 49   (100.0) 82.0±1.0 58   (100.0) 93.4±0.4
T=3,D=5 89   (100.0) 82.0±1.7 100   (100.0) 93.4±0.3
T=3,D=6 149   (100.0) 82.2±1.2 171   (100.0) 93.5±0.3
T=3,D=7 245   (100.0) 82.1±1.2 257   (100.0) 93.6±0.8
Appendix C Energy Efficiency Analysis
We utilize the maximum memory usage, inference latency and theoretical energy consumption as metrics to comprehensively evaluate the energy efficiency of various GT baselines. Specifically, we record the absolute elapsed running time per test epoch for GT-SVQ and other GT baselines. Notably, following the same settings as previous studies Wu et al. [2024], we use the mini-batch partition for training on the ogbn-products dataset.
The theoretical energy consumption estimation is derived from Yao et al. [2024]. In comparison, we fix some hyperparameters like the number of layers, the number of heads and the dimension of hidden embeddings for each model. The theoretical energy consumption of GTs during the inference phase is estimated in a straightforward way by counting floating point operations (FLOPs) and synaptic operations (SOPs). The energy cost of GT-SVQ can be formulated as follows:
$E$ $=\sum_{l=1}^{L}{(E_{SVQ}+E_{CGSA}+E_{MPNN}+E_{coding})+E_{CH}}$
$=\alpha_{s}(\sum_{l=1}^{L}\sum_{t=1}^{T}{SP}_{SVQ}^{l,t})+\alpha_{f}(\sum_{l=1}^{L}({FP}_{CGSA}^{l}+{FP}_{MPNN}^{l}$
$\qquad+{FP}_{coding}^{l})+{FP}_{CH})$ (25)
where $\alpha_{f}$ and $\alpha_{s}$, as scale factors for floating point and synaptic operations, are set to 4.5 and 0.9. $FP$ and $SP$ are denoted as floating point operations and synaptic operations of each layer. ${SP}^{l,t}=r^{l,t}\times{FLOP}^{l,t}$, where $r^{t,l}$ is the fire rate of spiking neurons in the $l$-th layer at the $t$-th time step. $E_{coding}$ derived from Zhou et al. [2022] is the energy consumption concerning the conversion between real-valued and spiking presentations in each layer. Notably, we can generate and store codewords corresponding to each node on the neuromorphic hardware. It enables GT-SVQ to maintain the codebook with relatively low energy consumption.
Figure 6: Accuracy versus Inference Time. The size of the circle indicates the maximum memory usage during model training.
The energy efficiency analysis in table 3 shows that GT-SVQ achieves the highest inference speedup on Physics. As depicted in Figure 6, GT-SVQ has achieved the highest accuracy (96.2%) and the fastest inference speed (21ms) among GT baselines on the Physics dataset. Compared with SVQ, the traditional strategy in VQ-based graph learning methods which measures distances and then replaces embeddings is inefficient. Besides, GT-SVQ achieves faster inference speed, better predictive performance and smaller memory footprint compared to another spike-based GT. We further construct the one-layer GT-SVQ. Benefiting from the fusion of SVQ and CGSA, GT-SVQ-1L can rapidly capture local and global graph information while bringing inference speed closer to that of standard GNNs like GAT.
Appendix D Revisting GT-SVQ in the perspective of homophily
There is a popular notion that message propagation-based methods are more suitable for graphs with high-level homophily Ma et al. [2021]. Therefore, in this section, we conduct a quantitative analysis to investigate whether the homophily of graphs is one of the determining factors on the performance of GT-SVQ. Specifically, we perform the same graph generation strategy on Cora and Citeseer datasets following the previous study Ma et al. [2021]. Figure 7 shows the influences of different homophily ratios on predictive results, where the homophily ratio is defined as |$\frac{|\{(v,w):(v,w)\in\mathcal{E}\cap y_{v}=y_{w}\}|}{|\mathcal{E}|}$. We find that as the homophily ratio decreases, the classification performance initially declines but eventually starts to improve. It is consistent with previous observations Ma et al. [2021] that the homophily assumption of message passing methods is not accurate. And it implies GT-SVQ may achieve strong performances on certain heterophilic graphs. The results in Table 2 show that GT-SVQ does achieve the best classification accuracy on the Actor dataset compared with other baselines. Empirical results highlight the efficacy of GT-SVQ on both heterophilic and homophilic graphs.
Figure 7: The accuracy of GT-SVQ on synthetic graphs (Cora and Citeseer) with various homophily ratios.
Appendix E Experimental Setup
Datasets.
We evaluate GT-SVQ on 9 datasets including three citation networks Sen et al. [2008] (Cora, CiteSeer, PubMed), two co-author networks Shchur et al. [2018] (Coauthor-CS and Coauthor-Physics ), two heterophily networks (Actor and Deezer) and two large-scale graphs (ogbn-arxiv and ogbn-products) from the Open Graph Benchmark (OGB) Hu et al. [2020]. For citation networks, the data splits adhere to the semi-supervised settings. For co-author networks and heterophily networks, we adopt the same splits in previous studies Li et al. [2023b]Wu et al. [2024]. For datasets from the OGB, we adopt their own standard splits.
Baselines.
To comprehensively evaluate the performance of GT-SVQ, a head-to-head comparison is conducted with state-of-the-art GNNs and GTs, based on their architectures. As shown in Table 1, components in baselines are fall into three categories: spike-based methods (SpikingGCN Zhu et al. [2022], SpikeNet Li et al. [2023a], SpikeGCL Li et al. [2023b], SpikeGT Sun et al. [2024]), Graph Transformer framework (NAGphormer Chen et al. [2022], NodeFormer Wu et al. [2022], SGFormer Wu et al. [2024], GOAT Kong et al. [2023]), vector quantization-based methods (VQGraph Yang et al. [2024]). Besides, three classic graph neural networks (GCN Kipf and Welling [2016], GAT Veličković et al. [2017], SGC Wu et al. [2019]) are also included in the comparison.
Appendix F Related Work
Spiking Neural Network.
Inspired by brain-like spiking computational frameworks, Spiking Neural Networks are proposed to address the computing energy consumption challenge. Different from Artificial Neural Networks (ANNs), neurons in SNNs communicate via binary and sparse spikes. It enables SNNs to reduce the storage overhead of intermediate outputs among layers and utilize more accumulation operations instead of multiply-accumulation operations Roy et al. [2019]. Some studies motivated by these advantages in energy efficiency have attempted to construct spike-driven neural networks. These networks can be broadly divided into two categories, ANN-to-SNN conversion and direct training framework. For the former, they tend to build a SNN upon a pre-trained ANN. These methods try to minimize information loss during the conversion process by performing scaling/normalizing operations on weights or replacing the activation layers with spiking neurons Diehl et al. [2015]Cao et al. [2015]Hao et al. [2023]. For the latter, studies try to directly train spike-driven neural networks by introducing surrogate gradient. This approach effectively reduces the strong dependence of SNNs on the number of time steps Fang et al. [2021a]Zheng et al. [2021].
Graph Transformers.
Although graph neural networks (GNNs) have become dominant paradigm cross various graph tasks Kipf and Welling [2016]Veličković et al. [2017], the message passing mechanism as the foundations of GNNs has some well-known drawbacks such as over-smoothing, over-squashing and the neglect of long-range information Li et al. [2018]Alon and Yahav [2020]. Graph Transformers (GTs), which can differentially aggregate messages over all nodes to alleviate local structure bias, have been developed to overcome above issues. Specifically, methods tend to inject the graph topology information into Transformer variants by introducing auxiliary GNNs or generating positional/structural embeddings from a graph. For some early studies, GTs are proposed to solve small-scale graph-level tasks like molecular property prediction and molecule classification Rampášek et al. [2022]Liao and Smidt [2022]. Recently, some methods aimed to enhance GTs’ performance on node-level tasks by constructing mini-batch sampling strategies and lightweight attention modules Shirzad et al. [2023]Li et al. [2024].
VQ-VAE and Follow-up.
To overcome the issue of posterior collapse, Van Den Oord et al. [2017] develop a discrete latent variational autoencoder (VAE) model called VQ-VAE. Input images are mapped to the embedding space through an encoder. The embeddings are replaced with the nearest pre-defined codebook entries through measuring distances between embeddings and entries. The replaced outputs are then fed into the decoder. It’s worth noting that VQ-VAE utilizes a commitment loss and the straight-through estimator to update the codebook and encoder-decoder modules. In addition, Kolesnikov et al. [2022] adopted a codebook splitting algorithm to improve codebook usage. Mentzer et al. [2023] implicitly constructs a finite scalar codebook by quantizing elements of intermediate embeddings to integers. In a nutshell, experiments have shown that VQ-VAE and its variants provide a simpler representation generation schema and an energy-efficient inference framework compared with their quantization-free counterparts.

### v2

@@ Introduction
where $L$ is the number of layers. We choose a simple fully connected layer as the classification head. The auxiliary MPNNs are implemented as a single-layer GCN without normalization. It has been observed that projection blocks, including Multilayer Perceptrons (MLPs) and normalization layers, exacerbate the overfitting problem on large-scale graphs for vanilla Transformers. Therefore, we discard projection layers and retain only the self-attention module and the residual connection (he2016deep).
Experiments
Comparison with Existing Models
In this section, we conduct the experimental evaluation to show the effectiveness of GT-SNT on nine node classification datasets, including three citation networks (kipf2016semi), two co-purchase networks (shchur2018pitfalls), two heterophilic graphs (wu2024sgformer) and two OGB graphs (hu2020open). A head-to-head comparison is conducted with state-of-the-art GNNs and GTs, based on their architectural designs. As shown in Table 1, these baselines fall into three categories: spike-based methods, Graph Transformer framework and vector quantization-based methods. The hyperparameter search strategy is deployed on both GT-SNT and other baselines to get the optimal combinations of parameters. We perform all models on each dataset 5 times with different random seeds to report the mean and standard deviation of metrics. All experiments are conducted on a single NVIDIA RTX 4090 GPU.
Overall performance.
The experimental results are demonstrated in Table 2. As shown in the table, our method achieves promising performance on all datasets. This is a significant advancement considering the information loss caused by low-precision spike count embeddings. Specifically, GT-SNT achieves predictive performances on par or even better than high-precision GT baselines. In general, our method shows slight improvements on 5 out of 9 datasets. GT-SNT effectively injects the graph inductive bias into the self-attention block while capturing long-range interactions, resulting in strong expressive power. Besides, we observe that GT-SNT has competitive performances on two heterophily datasets, particularly Actor. It suggests that GT-SNT can also maintain strong expressive power on heterophilic graphs. GT-SNT outperforms other spike-based baselines on all datasets, which achieves an average improvement of 3.9%. Constructing the input sequence from propagation steps rather than repeating graphs multiple times endows GT-SNT with desirable scalability. In addition, GT-SNT achieves a better trade-off between latent space compression and information retention by introducing spiking neurons as the node tokenizer instead of artificial neuron substitutes.
Models Components
SP GT VQ
SpikingGCN(zhu2022spiking) - -
SpikeNet(li2023scaling) - -
SpikeGCL(li2023graph) - -
SpikeGT(sun2024spikegraphormer) -
NAGphormer(chen2022nagphormer) - -
NodeFormer(wu2022nodeformer) - -
SGFormer(wu2024sgformer) - -
GOAT(kong2023goat) -
VQGraph(yang2024vqgraph) - -
GT-SNT
@@
#nodes\#nodes 2,708 3,327 19,717 18,333 34,493 7,600 28,281 169,343 2,449,029
#edges\#edges 10,556 9,104 88,648 163,788 495,924 30,019 185,504 1,166,243 61,859,140
@@
GT-SNT 84.7±0.8 74.0±0.5 80.6±0.4 93.7±0.4 96.2±0.0 39.1±0.2 65.7±0.1 72.4±0.3 74.8±0.4
Characteristics of Spiking Node Tokenization
To thoroughly analyze the spiking node tokenization, we conduct a series of experiments on the GT-SNT. We track the metrics including Codebook Usage and Accuracy to explore the following questions: (i) How does the size of the latent embedding space affect GT-SNT? (ii) Is the spiking node tokenization a more efficient tokenization alternative?
[…]
Figure 3: Influences of random feature dimensionality $D$ and propagation step $T$ on Cora and CS.
Figure 4: Codebook usage under four different latent space sizes on Cora and CS datasets. The solid-colored bar represents the number of used codewords, while the hatched bar denotes the size of the pre-defined codebook.
@@
For exploring the codebook collapse problem in VQ-based graph models, we track the fraction of codewords that are used at least once after the training phase. The codebook usages between GT-SNT and the other VQ-based graph methods are shown in Figure 4. In previous works, the pre-defined codebook is used to represent the entire discrete latent space. We select four combinations of $D$ and $T$ ($T=3/D=4$, $T=3/D=5$, $T=3/D=6$, $T=3/D=7$) to match the latent space sizes ($2^{8}$, $2^{10}$, $2^{12}$, $2^{14}$). The results demonstrate that those methods using pre-defined codebooks suffer from the serious issue of codebook collapse. As the pre-defined codebook size increases, the codebook usage decreases. For GOAT, the average codebook usages are 10.6% and 8.7% on Cora and CS datasets. The codebook usages in VQGraph are slightly higher, which achieve 16.9% and 29.0%. When the latent space size exceeds $2^{10}$, the usages of both methods drop below 50%. It reveals that these VQ-based graph learning methods struggle with efficiently using large pre-defined codebooks. SNT provides a generative solution, which dynamically selects the subset from a discrete spike count embedding space to reconstruct the codebook. It brings 100% codebook usage. Although the number of used codewords in GT-SNT is slightly larger than that of vanilla VQ counterparts in some cases, it is still significantly smaller than the pre-defined codebooks. In a nutshell, SNT achieves better codebook utilization by viewing spiking neurons as a sequence-to-token approach.
Datasets Metrics NAGphormer GOAT NodeFormer SGFormer SpikeGT Avg GT-SNT
@@
Energy Efficiency Analysis
To verify the efficiency of GT-SNT, we conduct energy efficiency analysis using the following metrics: peak memory usage during training, inference latency and theoretical energy consumption. The theoretical energy consumption estimation is derived from (yao2024spike). The results are presented in Table 3.
@@
The results show that GT-SNT achieves the lowest inference latency across all datasets. Compared to other GTs, GT-SNT with better performance achieves up to 130x lower inference latency. Most VQ-based methods tend to retrieve and replace node representations with the closest codewords during the inference phase. In SNT, trained spiking neurons directly convert input features into codewords, which reduces the computational cost brought by retrieval. Integrating the reconstructed codebook in SNT with linear-time global attention brings a significant improvement in inference speed. For large-scale graphs with high feature dimensionality, repeating the graph to generate sequential inputs leads to high energy consumption. The results show that previous spike-based GT is hard to manifest its superiority in energy consumption on CS and Physics. Through generating low-dimensional random features and collecting sequential embeddings from propagation steps, GT-SNT outperforms another spike-based GT in memory usage and latency across all datasets while maintaining acceptable theoretical energy consumption.
Ablation Study
@@
In this section, we conduct ablation studies to explore the influences of different components on predictive performances. We construct 8 baselines by combining two classic linear-time attention modules (Performer (choromanski2020rethinking) and Linformer (wang2020linformer)), two spiking neurons (IF and LIF), two normalization algorithms (LayerNorm (ba2016layernormalization) and STFNorm (xu2021exploiting)) and two VQ modules (VQ-VAE (van2017neural) and FSQ (mentzer2023finite)).
@@
The experimental results are demonstrated in Table 4. Although incorporating extra positional encodings enables Performer and Linformer to handle graph prediction tasks, they struggle to achieve good predictive performance on large-scale graphs like ogbn-arxiv. In GT-SNT, the CGSA actively introduces the global positional encoding during attention score calculation. It suggests that developing topology-aware tokenized Transformers is a promising direction for scaling GTs on large-scale graphs. The choice of spiking neurons also affects the predictive performances of GT-SNT. PLIF models with learnable membrane time constants and synaptic weights achieve better performances in most cases. These neurons effectively improve the flexibility of SNT. In addition, the well-designed normalization algorithm for spiking neurons, STFNorm outperforms the LayerNorm algorithm across all datasets. For complex tasks, GT-SNT with dynamic learning and adaptive codebook granularity surpasses the simple quantizers such as FSQ. Benefitting from diverse spiking dynamics and finer quantization, GT-SNT outperforms both FSQ and full-precision VQ-VAE.
@@
+FSQ 78.4±0.9 93.7±0.2 95.8±0.1 71.6±0.2
GT-SNT 80.6±0.5 93.7±0.4 96.2±0.0 72.4±0.3
Table 4: Ablation studies on PubMed, CS, Physics and ogbn-arxiv datasets. And +x+x means replacing the original component in GT-SNT with $x$.
Conclusion
@@
In this study, we propose GT-SNT, a linear-time Graph Transformer via spiking node tokenization. We find that positional encoding patterns of different nodes can be encoded into the same codewords. Inspired by the observation, GT-SNT not only considers spiking neurons as the low-power units, but also incorporates SNNs as a learnable tokenizer into Graph Transformers. It enables GT-SNT to achieve faster inference speed and better predictive performance. Spiking node tokenization that dynamically creates a spike count-based codebook paves a different way to address issues within current VQ-based graph models. We believe that our work holds great promise for the development of brain-inspired neural networks, which bridge the gap between existing SNNs and node tokenization. We hope it will inspire further research into energy-saving Graph Transformers.
Acknowledgments
This work was supported in part by the National Key R&D Program of China under Grant 2022YFF0902500, in part by the Guangdong Basic and Applied Basic Research Foundation, China under Grant 2023A1515011050, in part by Shenzhen Science and Technology Program under Grant KJZD20231023094501003, and in part by Tencent AI Lab under Grant RBFR2024004.